Aanyapie:

Two separate bacteria populations grow each month and are represented by the functions f(x) = 3x and g(x) = 7x + 6. In what month is the f(x) population greater than the g(x) population? (2 points) Month 1 Month 2 Month 3 Month 4

dude:

You probably meant an exponent for f(x)? \(f(x)=3^x\) and \(g(x)=7x+6\)

bammbiii5:

Well, you are given. \[f(x)=3^x \] and \[g(x)=7x+6\] So month 1 we can plug 1 in for x. So, x = 1 \[f(1)=3^1=3\] and \[g(1)=7*1+6=13\] So as you can tell: f(1) = 3 and g(1) = 13 so this is incorrect/ false. Because it is asking for f(x) to be larger not g(x). For B) month 2 plug 2 in for x just like we did above for month 1. \[f(2)=3^2=9 \] and \[g(2)=7*2+6=20\] So again as you can see: f(2) and g(2) So this is incorrect/ false. g(x) is larger than f(x). For C) Month three plug 3 in for x again like we did above. ^^ \[f(3)=3^3=27\] and \[g(3)=7*3+6=27 \] So this is incorrect because both are equal. f(x) = g(x) For D) Month 4. Plug 4 in for x like above ^^. \[f(4)=3^4=81\] and \[g(4)=7*4+6=34\] This one is correct because f(x) > g(x) or f(x) is greater than g(x). Answer is D) Month 4. Good luck on the assignment.